1. What is the oxidation state of copper in CuSO3 is: a.2 b.–1 c.–2 d.1 e.0

2. A solution where [H+] = 10-13 M is ______________. a.acidic b.two of these c.neutral d.strongly acidic e.basic

3. Consider the following reaction: Mg + Br2 → MgBr2

Which of the following statements is false?

a.The bromine atom is losing electrons; therefore, it is reduced.

b.The bromine atom is gaining electrons; therefore, it is oxidized.

c.The magnesium atom is losing electrons; therefore, it is oxidized.

d.none of these

e.The magnesium atom is gaining electrons; therefore, it is oxidized

4. 8.55 x 10-3 moles of an unknown compound has a mass of 0.257 g. The compound could be: Select one: a. C2H6 b. CH4 c. C5H12 d. C3H8   e. C4H10

5.ΔH for the reaction below is -826.0 kJ/mol. Calculate the heat change when a 69.03-g sample of iron is reacted. 4 Fe(s) + 3 O2(g) → Fe2O3(s)

Select one: a. –2.851 × 104 kJ b. –1020.9 kJ c. –2042 kJ d. –255.2 kJ e. –510.5 kJ

6. An example of a secondary structure of a protein is

Select one: a. a pleated sheet b. none of these c. an alpha-amino acid d. a peptide linkage e. serine

7. For a double replacement reaction to occur, one of the products must be insoluble, determine which reaction(s) will NOT occur using your solubility tables. Select one or more:

a. AgNO3+ MgCl2 →

b. Al(NO3)3 + NaOH →

c. NaNO3 + KI →

d. Pb(NO3)2 + KI →

e. CuBr2 + AlCl3 →

(d is the right choice in Q.7 can you select any other which is the right choice apart from d)
Answer #1

Ans 1

The given compound = CuSO3

Oxidation number of S = +6

Oxidation number of O = -2

Oxidation number of Cu = x

the sum of all oxidation numbers must be zero.

x + (+6) + 3*(-2) = 0

x = - 6 + 6 = 0

Option E is the correct answer